OpenStudy (anonymous):

What is √180 rounded to the nearest hundredth?Select one of the options below as your answer: A. 13.41 B. 13.42 C. 13.43 D. 13.44

OpenStudy (radar):

\[\sqrt{36}\sqrt{5}=6\sqrt{5}13.42\]B.

OpenStudy (anonymous):

easy The answer is A.

OpenStudy (anonymous):

no... 13.41640786 is rounded to 13.41

OpenStudy (radar):

Insert equal sign after sq rt 5

OpenStudy (anonymous):

that doesnt make sense to me still

OpenStudy (anonymous):

the question is straightforward, u dont need all that work

OpenStudy (radar):

\[6\sqrt{5}=13.42\]When you have memorizes the square root of 2, 3, 5, and 7 it does make sense lol

OpenStudy (anonymous):

Please listen to me person. I am sorry, but michael jordan listen to me

OpenStudy (radar):

Also the result is closer to 13.42 than 13.41

OpenStudy (anonymous):

AAAH

OpenStudy (anonymous):

i didnt know 13. 42 was a choice. Sorry

OpenStudy (radar):

The results is 13.4164 Note the 6 is greather than 5 so round up.

OpenStudy (anonymous):

I noticed again. Thank you for sir.

OpenStudy (radar):

I am surprised that the answers are so close. Thanks for your input

OpenStudy (anonymous):

me too

OpenStudy (radar):

Good Day, and enjoy "Open Study"

OpenStudy (anonymous):

square root(180)=13.4164079 rounded to the nearest hundreth Answer-->13.42